an experienced painter can paint a garage twice as fast as his apprentice. working together, the painters require 4 hours to paint a garage. how long would it take the experienced painter working alone to paint the garage?
Answer by rwm(914)

4/x+4/2x=1
4/x+2/x=1
6/x=1
x=6 hrs for the experienced painter
2x=12 for the apprentice.